### v1

@@ Abstract
A unified framework is provided for analysing check-all-that-apply (CATA) product data following the “one citation, one vote” principle. CATA data arise from studies where A consumers evaluate P products by describing samples by checking all of the T terms that apply. Giving every citation the same weight, regardless of the assessor, product, or term, leads to analyses based on the L1 norm where the median absolute deviation is the measure of dispersion. Five permutation tests are proposed to answer the following questions. Do any products differ? For which terms do products differ? Within each of the terms, which products differ? Which product pairs differ? On which terms does each product pair differ? Additionally, we show how products and terms can be clustered following the “one citation, one vote” principle and how L1-norm principal component analysis (L1-norm PCA) can be applied to visualize CATA results in few dimensions. Together, the permutation tests, clustering methods, and L1-norm PCA provide a unified approach. The proposed methods are illustrated using a data set in which 100 consumers evaluated 11 products using 34 CATA terms. R code is provided to perform the analyses.
@@
check-all-that-apply (CATA), pick-any, permutation tests, median absolute deviation (MAD), L1 norm, L1 norm principal component analysis (L1-norm PCA)
@@ 1 Introduction
Sensory and consumer studies are often designed to investigate how humans perceive foods, beverages, or non-food products. These studies usually involve structured sample evaluations of the products by human assessors with data analysis by sensometric methods. Some studies obtain precise sensory characterizations of the products from panelists trained to identify and scale sensory attributes from a consensus lexicon. Other studies investigate how consumers perceive and conceptualise products. Both types of studies frequently use the check-all-that-apply (CATA) question format (Meyners and Castura, 2014; Ares and Jaeger, 2023), which provides a list of terms to allow assessors to characterise products. For example, a CATA question might instruct the assessor to check as many sensory or emotion terms from the list as describe the product sample under evaluation. Such a study leads to a three-way array (assessors-by-products-by-terms) of raw data.
@@
Since checked terms, called citations or elicitations, can be analysed as count data, CATA data are often aggregated over the assessors to obtain a two-way products-by-terms table of frequencies. Thus, a CATA table counts the assessors who cited each term for each product.
@@
CATA frequency tables are often analysed as contingency tables (Dooley et al, 2010; Valentin et al, 2012; Meyners et al, 2013; Meyners and Castura, 2014), in which a normalisation or adjustment is applied that in some sense equalises the contributions of variables with different citation rates. The widely used normalisation inherent in the chi-squared test has this feature: each cell frequency is expressed relative to the reciprocal of the square root of its expectation, given the column and row marginal sums. However, a CATA frequency table is not a contingency table. In a contingency table, each member of the assessor sample contributes to only one table cell. In a CATA frequency table, each cell indicates how many assessors cited a particular term (column) for a particular product (row). For this reason, we propose to analyse raw counts without normalising terms or products. We will investigate differences between products for each term separately (univariately) and over all terms simultaneously (multivariately), using the novel approach of treating each citation equally. Our proposal is embodied in the following principle: one citation, one vote!
@@
This proposal provides a new way of summarising and analysing CATA data that we have not seen used previously. Treating every citation equally leads us to use medians (not means) as our measure of central tendency, medians of absolute deviations from medians as our measure of dispersion (not variances), and median absolute deviations (not Euclidean distances) to quantify differences when grouping products and grouping terms in cluster analysis. Additionally, we use L1-norm principal component analysis (L1-norm PCA; Ke and Kanade, 2005), rather than conventional PCA, to summarise the product information visually in few dimensions.
@@
Data and methods are introduced in Section 2. The data come from a study involving 100 consumers who used 34 emotion terms to evaluate 11 blackcurrant squash products (fruity beverages) available commercially in the United Kingdom (Ng et al, 2013a). There, we also describe the proposed statistical methods, all of which are aligned with the “one citation, one vote” principle. We describe how to conduct distribution-free permutation tests, as well as how to adjust for multiple testing. We describe how to perform cluster analysis of products and terms, as well as how to explore relationships between products and terms using the L1-norm PCA. Section 3 presents the results obtained from applying these methods to the blackcurrant squash CATA data. We discuss these results and make conclusions in Section 4.
@@ 2.1 Data from the blackcurrant squash study
This paper uses data extracted from research by Ng et al (2013a) in which consumers evaluated 11 commercial blackcurrant squashes, representing a range of sensory and packaging properties observed in the UK market. The original labelling of the 11 squash products (P1 to P11) in Ng et al (2013a) is supplemented here by two additional letters (ab). Letter a indicates the economical, niche and standard market-value categories, respectively a = e, n, s. Letter b indicates products with and without added sugar, respectively b = w, o. For example, the first product P1sw is a standard product with added sugar.
A preliminary study described by Ng et al (2013b) was conducted to select relevant conceptual terms with emotional or functional connotations (Thomson et al, 2010) for evaluating blackcurrant squashes under three conditions: blind, pack, and informed. In this manuscript, we analyse results from the blind condition only, in which consumers evaluated each sample without seeing the package. In this condition, consumers’ emotional conceptualizations of the samples are not influenced by information about brand, assumptions about the market-value category (niche vs standard vs economical), added sugar, or use of non-nutritive sweeteners, if any.
@@
The three-way raw data, depicted schematically in Figure 1 contains indicator variables: zeros indicate terms not cited; ones indicate terms cited. From these raw data, the CATA frequency table is the $P\times T$ table formed by summing the $A$ slices across the assessors.
@@ 2.2 A new approach, simple and robust
To treat every response equally, we will measure differences by their absolute values and use the median as a measure of centrality. The median is the value that minimises the sum of absolute differences to the sample values (Schwertman et al, 1990). We measure dispersion about the median by finding the median of the absolute deviations between the sample values and the median, called the median absolute deviation (MAD) about the median. Since absolute differences are less sensitive to outliers, the median and MAD are more robust measures of central tendency and overall difference, and fit the “one citation, one vote” principle.
@@
We draw attention to three points. First, as just illustrated, the measures of central tendency and dispersion usually increase in tandem. When least-squares methods are applied in multivariate analyses, terms having unequal variances are often rescaled to equalise their contributions to the analysis; for example, in correspondence analysis (Benzécri, 1973; Greenacre, 2016), the chi-square normalisation rescales variables and samples so all are on the same footing. In our case, following the “one citation, one vote” principle, we do not attempt to balance the influence of any terms, products, or assessors. Larger product differences between products based on citations of the term Happy are considered to discriminate the products better than the smaller differences between products based on citations of the term Sickly. Second, the summary values used (median, MAD) are very easy to understand and interpret because they are measured in percentages, which is exactly the same scale used in the CATA table. Third, the median and the MAD are more robust to outliers than measures based on squared differences, such as the mean and variance, which are influenced more strongly by outlying sample values.
@@ 2.3 Permutation testing
To test differences between products on the selected terms, we make use of the simple, non-parametric approach called permutation testing; for a complete treatment, see Good (2004), and for previous applications in the analysis of CATA tables, see Meyners et al (2013) and Mahieu et al (2021). As we will show, this approach can be used to test various hypotheses, such as whether individual cells in the CATA table are larger or smaller than expected or whether a pair of products differs on any of the terms. The approach naturally generalises to multivariate tests of whether products differ globally, accounting for all terms simultaneously.
@@
Now, we describe the procedure in more specific detail. Operationally, each simulation takes the form of random permutations of the product labels within each assessor. Figure 1 shows the set of citations in the $P\times T$ matrix “slice” of the array corresponding to a particular assessor. Permuting the $P$ rows of this matrix assigns rows to different product labels but preserves dependencies in elicitations across the multivariate terms. For example, a particular assessor who tends to endorse the terms Pleased and At ease together or not at all has data dependencies that are kept intact by permuting the rows (products) keeping the columns (terms) intact. This procedure is applied to data from all $A$ assessors. Since rows are assigned to product labels independently per assessor, any two assessors will usually not have the same product rows assigned to the same product labels. After permuting product rows within assessors, assessor matrix slices are aggregated into a simulated $P\times T$ CATA table. Repeating this procedure $B$ times yields $B$ random CATA tables, which together with the original CATA table, gives $B+1$ CATA tables in total.
@@ 2.4 Multiple testing and false discovery
Doing the pairwise test for each term $t$ (Test 5), the null distribution is comprised of the test statistic, which is the observed signed difference observed, along with the similar differences in the $B$ random CATA tables. The p-value is again the proportion of absolute outcomes in the null distribution as large or larger than the absolute test statistic. It is possible to test only product pairs determined to differ significantly based on the multivariate test (Test 4). Here, all product pairs on all terms are tested using the BH procedure, and the results are visualised for each term in a $P\times P$ heatmap.
@@ 2.5 Hierarchical cluster analysis
2.6 L1-norm principal component analysis (L1-norm PCA)
Without loss of generality, suppose matrix $\bf X$ has median-centred columns. L1-norm PCA decomposes $\bf X$ into two matrices: a score matrix $\bf T$ and loading matrix $\bf P$. Dimension reduction is possible since retaining only $K$ components minimises the sum of absolute residuals
@@ 2.6 L1-norm principal component analysis (L1-norm PCA)
where T denotes matrix transposition and $\|\cdot\|_{1}$ denotes the L1-norm, that is, the sum of absolute elements. Several L1-norm PCA algorithms have been proposed for approximating ${\bf X}$ with ${\bf T}_{K}{\bf P}_{K}^{\sf T}$. We use the algorithm proposed by Ke and Kanade (2005).
Additionally, we show the uncertainty of the products in the L1-norm PCA visualization. To do so, we conduct bootstrapping at the assessor level. Each assessor has $PT$ values in his or her set of responses (a slice of the array shown in Figure 1.) Assessors are randomly chosen with replacement, until another three-way array of $A$ slices, as shown in Figure 1, is constructed, where some assessors are included more than once, some not at all. From this bootstrapped array another CATA table is formed, and its rows are projected onto the original solution already established. After repeating this bootstrap procedure 1000 times, each time projecting the rows to the solution, the points are overlaid with 95% covering ellipses, which estimate the confidence regions for the products. If confidence ellipses on two product means do not overlap, this is an indication, but not conclusive proof, of a significant difference between the products. In the present L1-norm framework, we prefer to judge this pairwise product difference using the permutation test previously proposed.
In order to quantify the success of the L1-norm PCA in visualising the matrix $\bf X$, the sum of the absolute residuals in Eq. (2) can be expressed relative to $\|{\bf X}\|_{1}$ and subtracted from 1, to give the proportion explained by the $K$-dimensional L1 approximation:
$\textsf{prop}_{K}=1-\frac{\|{\bf X}-{\bf T}_{K}{\bf P}_{K}^{\sf T}\|_{1}}{\|{\bf X}\|_{1}}$ (3)
The solutions in L1-norm PCA are not nested as in regular PCA, so percentages of explanation cannot be ascribed to individual dimensions, only to specific $K$-dimensional solutions. To obtain a scree plot similar to that usually produced in PCA, the L1-norm PCA is performed separately for solutions of dimension $K=1,2,\ldots,{\rm min}\{P,T\}$, and the gains in explanation are then computed, going from the solutions of dimension $K-1$ to $K$ (where, for $K=0$, the explanation is 0).
@@ 2.7 Software
Data analyses were conducted in R 4.2.2 (R Core Team, 2024). Functions madperm() and mad.dist() in the R package cata (Castura, 2025) were used to conduct permutation tests and to calculate MAD-based distances for cluster analyses. R code in Supplementary Material shows how to reproduce the permutation tests presented here. Functions l1pca() and l1projection() in the R package pcaL1 (Jot et al, 2023) were used to conduct L1-norm PCA and to project bootstrapped sample points onto the L1-norm PCA biplots. Function CIplot_biv() in the R package easyCODA (Greenacre, 2018) used these points to calculate and draw confidence ellipses. Heatmaps and dendrograms were drawn using the function pheatmap() in the R package pheatmap (Kolde, 2019).
@@ 3.1 Tests and heatmaps
Test 1 conducts a global multivariate test to determine whether product differences exist on any term. The test statistic is the median of the MAD values from the $34$ terms, and is equal to 5%. Figure 2 shows the null distribution for this test statistic.
Figure 2: Null distribution values for the overall test statistic for product differences (x axis) vs proportion of occurrence (y axis). The median absolute deviance (MAD) is calculated for each term; the test statistic is the median of the MAD values across the 34 terms, equal to 5%.
@@
Next, we investigate which of the $T$ terms discriminate the products. For the term Happy, citation percentages range from 15% to 55%. The MAD value for Happy is 10%, which is the median “distance” of the 11 citation percentages from the median of 31%, on the percentage scale. Figure 3 shows graphically both the median (as a wide vertical bar) and MAD (as a narrow bar on top) for each term. We emphasise the MAD bars of terms that significantly discriminate the products, as determined by the univariate test results (Test 2), which are described next.
@@
Table 2: Results for the univariate test (Test 2) controlling the false discovery rate controlled at 0.05 show sorted p-values alongside the Benjamini-Hochberg (BH) series. The 22 terms having a p-value below the critical value (0.0125) were considered significant. The non-significant terms below the double line did not discriminate the products.
Figure 4 (top row) shows a heatmap visualisation of the univariate (Test 2) p-values in Table 2. Terms that significantly discriminate the products are shown in green, with deeper colours indicating a more significant test result. Terms above the BH critical value, which are non-significant across the products, are shown in light grey.
@@
Cluster analysis of the terms, as described in Section 2.5, organises the terms into a two-cluster solution, shown at the top of Figure 4. Positively valenced terms (cluster 1) are contrasted against negatively valenced terms (cluster 2). Term cluster 1 consists of 15 positively valenced terms, 11 of which discriminated the products: Comforted, Guilty pleasure, Trust, Approval, Pleasant surprise, Interested, At ease, Pleased, Satisfaction, Happy and Good. This grouping suggests consumers might have conceptualised Guilty pleasure as positively valenced, but the lexicon developers did not classify it either way, so it was considered unclassified (Ng et al, 2013a). The remaining four terms in this cluster (Desire, Warm, Attentive, Reminiscence) did not discriminate the products. Term cluster 2 is comprised of 19 negatively valenced terms, of which 11 discriminated the products significantly (Disappointment, Displeasure, Annoyed, Shocked, Discontent, Uncomfortable, Disgust, Unhappy, Unpleasant surprise, Sickly, Regret) and eight did not (Bored, Curious, Angry, Worried, Cautious, Confused, Sceptical, Resentment).
Figure 4 (lower part) also shows a $P\times T$ heatmap visualization of the elementwise (Test 3) p-values obtained by analysing, for each term, which of the $P$ products differs from the median value. We decided a priori to conduct all $P\times T=11\times 34=374$ elementwise tests (Section 2.3) with the FDR controlled at level 0.05 (Section 2.4). In 374 elementwise tests, 89 p-values were less than or equal to the BH critical value of $0.0118$.
@@
At least one product differed significantly from the median for all but four terms: Attentive, Worried, Cautious, and Confused. These four terms are among the 12 non-significant terms (coloured light grey in the top row of Figure 4). Inspection of the columns for the other eight non-significant terms show at least one significant elementwise test. This result shows that individual products can differ significantly from the median of a term, even if the product citation percentages do not differ significantly overall on that term.
@@
Figure 4 shows the dendrogram from the complete-linkage cluster analysis of products to the left of the elementwise heatmap. The rows (products) are sorted to accommodate this dendrogram which shows a crisp two-cluster solution. Product cluster 1 is composed of five standard products with added sugar (P5sw, P7sw, P11sw, P1sw, P9sw). Product cluster 2 is composed of six more varied products: four products without added sugar (P10eo, P8so, P2so and P6so) and two niche products with added sugar (P3nw and P4nw).
@@
The elementwise heatmap reveals approximate characterisations of products in the two product clusters. Members of product cluster 1 elicited significantly higher citation percentages than the median product in positively valenced terms in term cluster 1 and significantly lower citation percentages than the median product in negatively valenced terms in term cluster 2. By contrast, members of product cluster 2 elicited significantly higher citation percentages than the median product in negatively valenced terms in term cluster 2 and significantly higher citation percentages than the median product in positively valenced terms in term cluster 1.
@@
We draw attention to selected product results (rows) in the elementwise heatmap related to product cluster 2. Like other members of product cluster 2, the economical product P10eo had higher-than-median citation percentages for negatively valenced terms in term cluster 2 and lower-than-median citation percentages for positively valenced terms in term cluster 1. Additionally, P10eo elicited significantly higher-than-median citation percentages for Bored, Angry, Sceptical and Resentment and a significantly lower-than-median citation percentage for Curious. Unlike the other 10 products, P10eo had relatively high intensities of catty, confectionery blackcurrant and artificial sweetness (Ng et al, 2012), which might explain why this product elicited these negatively valenced terms.
@@
Three standard products without added sugar (P2so, P6so and P8so) were also allocated to cluster 2. Besides of the general trend of product cluster 2, P6so and P8so elicited higher-than-median citation percentages for Sickly in term cluster 2. In a previous study, P6so and P8so were observed to have the highest intensities of artificial sweetness and natural processed blackcurrant (Ng et al, 2012), suggesting the relatively high Sickly citation rates were due to cloying oversweetness and processed flavour. These same sensory properties could be the reason for a lower-than-median citation percentage of Trust for P6so.

### v2

@@ Abstract
A unified framework is provided for analysing check-all-that-apply (CATA) product data following the “one citation, one vote” principle. CATA data arise from studies where A assessors evaluate P products by describing samples by checking all of the T terms that apply. Giving every citation the same weight, regardless of the assessor, product, or term, leads to analyses based on the L1 norm where the median absolute deviation is the measure of dispersion. Five permutation tests are proposed to answer the following questions. Do any products differ? For which terms do products differ? Within each of the terms, which products differ? Which product pairs differ? On which terms does each product pair differ? Additionally, we show how products and terms can be clustered following the “one citation, one vote” principle and how principal component analysis using the L1-norm (L1-PCA) can be applied to visualize CATA results in few dimensions. Together, the permutation tests, clustering methods, and L1-PCA provide a unified approach that provides robust results measured in citation percentages. The proposed methods are illustrated using a data set in which 100 consumers evaluated 11 products using 34 CATA terms. R code is provided to perform the analyses.
@@
check-all-that-apply (CATA), pick-any, permutation tests, median absolute deviation (MAD), L1 norm, L1-norm principal component analysis (L1-PCA)
@@ 1 Introduction
Sensory and consumer studies are often designed to investigate how humans perceive foods, beverages, or non-food products. These studies usually involve structured sample evaluations of the products by human assessors with data analysis by sensometric methods. Some studies obtain precise sensory characterizations of the products from panelists trained to identify and scale sensory attributes from a consensus lexicon. Other studies investigate how consumers perceive and conceptualise products. Both types of studies frequently use the check-all-that-apply (CATA) question format (Ares & Jaeger, 2023; Meyners & Castura, 2014), which provides a list of terms to allow assessors to characterise products. For example, a CATA question might instruct the assessor to check as many sensory or emotion terms from the list as describe the product sample under evaluation. Such a study leads to a three-way array (assessors-by-products-by-terms) of raw data.
@@
Since checked terms, called citations or elicitations, can be analysed as count data, CATA data are often aggregated over the assessors to obtain a two-way products-by-terms table of frequencies. Thus, a CATA frequency table counts the assessors who cited each term for each product.
@@
CATA frequency tables are often analysed as contingency tables (Dooley et al., 2010; Mahieu et al., 2021; Meyners & Castura, 2014; Meyners et al., 2013; Valentin et al., 2012), in which a normalisation or adjustment is applied that in some sense equalises the contributions of variables with different citation rates. The widely used normalisation inherent in the chi-squared test has this feature: each cell frequency is expressed relative to the reciprocal of the square root of its expectation, given the column and row marginal sums. However, a CATA frequency table is not a contingency table. In a contingency table, each member of the assessor sample contributes to only one table cell. In a CATA frequency table, each cell indicates how many assessors cited a particular term (column) for a particular product (row). For this reason, we propose to analyse raw counts without normalising terms or products. We will investigate differences between products for each term separately (univariately) and over all terms simultaneously (multivariately), using the novel approach of treating each citation equally. Our proposal is embodied in the following principle: one citation, one vote!
@@
This proposal provides a new way of summarising and analysing CATA data that we have not seen used previously. Treating every citation equally leads us to use medians (not means) as our measure of central tendency, medians of absolute deviations from medians as our measure of dispersion (not variances), and median absolute deviations (not Euclidean distances) to quantify differences when grouping products and grouping terms in cluster analysis. Additionally, we use L1-norm principal component analysis (L1-PCA; Ke & Kanade, 2005), rather than conventional L2-norm PCA (L2-PCA), to summarise the product information visually in few dimensions.
@@
Data and methods are introduced in Section 2. The data come from a study involving 100 consumers who used 34 emotion terms to evaluate 11 blackcurrant squash products (fruity beverages) available commercially in the United Kingdom (Ng et al., 2013a). There, we also describe the proposed statistical methods, all of which are aligned with the “one citation, one vote” principle. We describe how to conduct distribution-free permutation tests, as well as how to adjust for multiple testing. We describe how to perform cluster analysis of products and terms, as well as how to explore relationships between products and terms using L1-PCA. Section 3 presents the results obtained from applying these methods to the blackcurrant squash CATA data. We discuss these results and make conclusions in Section 4.
@@ 2.1 Data from the blackcurrant squash study
This paper uses data extracted from research by Ng et al. (2013a) in which consumers evaluated 11 commercial blackcurrant squashes, representing a range of sensory and packaging properties observed in the UK market. The original labelling of the 11 squash products (P1 to P11) in Ng et al. (2013a) is supplemented here by two additional letters (ab). Letter a indicates the economical, niche and standard market-value categories, respectively a = e, n, s. Letter b indicates products with and without added sugar, respectively b = w, o. For example, the first product P1sw is a standard product with added sugar.
A preliminary study described by Ng et al. (2013b) was conducted to select relevant conceptual terms with emotional or functional connotations (Thomson et al., 2010) for evaluating blackcurrant squashes under three conditions: blind, pack, and informed. In this manuscript, we analyse results from the blind condition only, in which consumers evaluated each sample without seeing the package. In this condition, consumers’ emotional conceptualizations of the samples are not influenced by information about brand, assumptions about the market-value category (niche vs standard vs economical), added sugar, or use of non-nutritive sweeteners, if any.
@@
The three-way raw data, depicted schematically in Fig. 1 contains indicator variables: zeros indicate terms not cited; ones indicate terms cited. From these raw data, the CATA frequency table is the $P\times T$ table formed by summing the $A$ slices across the assessors.
@@ 2.2 A new approach, simple and robust
To treat every response equally, we will measure differences by their absolute values and use the median as a measure of centrality. The median is the value that minimises the sum of absolute differences to the sample values (Schwertman et al., 1990). We measure dispersion about the median by finding the median of the absolute deviations between the sample values and the median, called the median absolute deviation (MAD) about the median. Since absolute differences are less sensitive to outliers, the median and MAD are more robust measures of central tendency and overall difference, and fit the “one citation, one vote” principle.
@@
We draw attention to three points. First, as just illustrated, the measures of central tendency and dispersion usually increase in tandem. When least-squares methods are applied in multivariate analyses, terms having unequal variances are often rescaled to equalise their contributions to the analysis; for example, in correspondence analysis (Benzécri, 1973; Greenacre, 2016), the chi-square normalisation rescales variables and samples so all are on the same footing. In our case, following the “one citation, one vote” principle, we do not attempt to balance the influence of any terms, products, or assessors. Larger differences between products based on citations of the term Happy are considered to discriminate the products better than the smaller differences between products based on citations of the term Sickly. Second, the summary values used (median, MAD) are very easy to understand and interpret because they are measured in percentages, which is exactly the same scale used in the CATA table. Third, the median and the MAD are more robust to outliers than measures based on squared differences, such as the mean and variance, which are influenced more strongly by outlying sample values.
@@ 2.3 Permutation testing
To test differences between products on the selected terms, we make use of the simple, non-parametric approach called permutation testing; for a complete treatment, see Good (2004), and for previous applications in the analysis of CATA tables, see Meyners et al. (2013) and Mahieu et al. (2021). As we will show, this approach can be used to test various hypotheses, such as whether individual cells in the CATA table are larger or smaller than expected or whether a pair of products differs on any of the terms. The approach naturally generalises to multivariate tests of whether products differ globally, accounting for all terms simultaneously.
@@
Now, we describe the procedure in more specific detail. Operationally, each simulation takes the form of random permutations of the product labels within each assessor. Fig. 1 shows the set of citations in the $P\times T$ matrix “slice” of the array corresponding to a particular assessor. Permuting the $P$ rows of this matrix assigns rows to different product labels but preserves dependencies in elicitations across the multivariate terms. For example, a particular assessor who tends to endorse the terms Pleased and At ease together or not at all has data dependencies that are kept intact by permuting the rows (products) keeping the columns (terms) intact. This procedure is applied to data from all $A$ assessors. Since rows are assigned to product labels independently per assessor, any two assessors will usually not have the same product rows assigned to the same product labels. After permuting product rows within assessors, assessor matrix slices are aggregated into a simulated $P\times T$ CATA table. Repeating this procedure $B$ times yields $B$ random CATA tables, which together with the original CATA table, gives $B+1$ CATA tables in total.
@@ 2.4 Multiple testing and false discovery
Doing the pairwise test for each term $t$ (Test 5), the null distribution is comprised of the test statistic, which is the observed signed difference, along with the similar differences in the $B$ random CATA tables. The p-value is again the proportion of absolute outcomes in the null distribution as large or larger than the absolute test statistic. It is possible to test only product pairs determined to differ significantly based on the multivariate test (Test 4). Here, all product pairs on all terms are tested using the BH procedure, and the results are visualised for each term in a $P\times P$ heatmap.
@@ 2.5 Hierarchical cluster analysis
2.6 L1-norm principal component analysis (L1-PCA)
Without loss of generality, suppose matrix $\bf X$ has median-centred columns. L1-PCA decomposes $\bf X$ into two matrices: a score matrix $\bf T$ and loading matrix $\bf P$. Dimension reduction is possible since retaining only $K$ components minimises the sum of absolute residuals
@@ 2.6 L1-norm principal component analysis (L1-PCA)
where T denotes matrix transposition and $\|\cdot\|_{1}$ denotes the L1 norm, that is, the sum of absolute elements. Several L1-PCA algorithms have been proposed for approximating ${\bf X}$ with ${\bf T}_{K}{\bf P}_{K}^{\sf T}$. We use the algorithm proposed by Ke & Kanade (2005).
Additionally, we show the uncertainty of the products in the L1-PCA visualization. To do so, we conduct bootstrapping at the assessor level. Each assessor has $PT$ values in his or her set of responses (a slice of the array shown in Fig. 1.) Assessors are randomly chosen with replacement, until another three-way array of $A$ slices, as shown in Fig. 1, is constructed, where some assessors are included more than once, some not at all. From this bootstrapped array another CATA table is formed, and its rows are projected onto the original solution already established using linear programming to minimize the sum of absolute deviances, subject to constraints (Ke & Kanade, 2005). The L1-PCA algorithm we used (Ke & Kanade, 2005) is robust to initialization values and reaches more precise solutions by allowing more iterations towards a smaller convergence tolerance. A solution could be considered converged if changing the initialization values, allowing more iterations, and specifying a smaller convergence tolerance does not provide meaningful improvement. After repeating this bootstrap procedure 1000 times, each time projecting the rows to the solution, the points are overlaid with 95% covering ellipses, which estimate the confidence regions for the products. If confidence ellipses on two product means do not overlap, this is an indication, but not conclusive proof, of a significant difference between the products. In the present L1 norm framework, we prefer to judge this pairwise product difference using the permutation test previously proposed.
The total dispersion is the L1 norm of $\bf X$. The L1-PCA with $K$ components extracts the largest possible sum of absolute deviances from $\bf X$. The proportion of the L1 norm extracted in these $K$ components,
$\textsf{prop}_{K}=\frac{\|{\bf T}_{K}{\bf P}_{K}^{\sf T}\|_{1}}{\|{\bf X}\|_{1}}$ (3)
quantifies the success of this L1-PCA in approximating $\bf X$. The solutions in L1-PCA are not nested as in regular PCA, so percentages of explanation cannot be ascribed to individual dimensions, only to specific $K$-dimensional solutions. To obtain a scree plot similar to that usually produced in PCA, the L1-PCA is performed separately for solutions of dimension $K=1,2,\ldots,{\rm min}\{P,T\}$, and the gains in explanation are then computed, going from the solutions of dimension $K-1$ to $K$ (where, for $K=0$, the explanation is 0). Different researchers will make different choices for how many components to retain. Since we were interested in product separation, we kept only the first $K$ components that separated products, i.e., where one or more product confidence ellipse excluded the origin.
@@ 2.7 Software
Data analyses were conducted in R 4.2.2 (R Core Team, 2024). Functions madperm() and mad.dist() in the R package cata (Castura, 2025) were used to conduct permutation tests and to calculate MAD-based distances for cluster analyses. R code in Supplementary Material shows how to reproduce the permutation tests presented here. Running these permutation tests took approximately about four minutes on a laptop running 64-bit Windows 10 Pro 22H2 with an Intel® Core™ i5-1235U, 1300 Mhz, 10 Core(s), 12 Logical Processor(s) and 16 GB RAM. Functions l1pca() and l1projection() in the R package pcaL1 (Jot et al., 2023) were used to conduct L1-PCA and to project bootstrapped sample points onto the L1-PCA biplots. Function CIplot_biv() in the R package easyCODA (Greenacre, 2018) used these points to calculate and draw confidence ellipses. Heatmaps and dendrograms were drawn using the function pheatmap() in the R package pheatmap (Kolde, 2019).
@@ 3.1 Tests and heatmaps
Test 1 conducts a global multivariate test to determine whether product differences exist on any term. The test statistic is the median of the MAD values from the $34$ terms, which was 5% (at the extreme right).. Fig. 2 shows the null distribution for this test statistic.
Figure 2: Null distribution values for the overall test statistic for product differences (x axis) vs proportion of occurrence (y axis). The median absolute deviance (MAD) is calculated for each term; the observed test statistic is the median of the MAD values across the 34 terms, which was 5% (at the extreme right).
@@
Next, we investigate which of the $T$ terms discriminate the products. For the term Happy, citation percentages range from 15% to 55%. The MAD value for Happy is 10%, which is the median “distance” of the 11 citation percentages from the median of 31%, on the percentage scale. Fig. 3 shows graphically both the median (as a wide vertical bar) and MAD (as a narrow bar on top) for each term. This plot shows the median and MAD values are correlated. We emphasise the MAD bars of terms that significantly discriminate the products, as determined by the univariate test results (Test 2), which are described next.
@@
Table 2: Results for the univariate test (Test 2) controlling the false discovery rate at 0.05 show sorted p-values alongside the Benjamini-Hochberg (BH) series. The 22 terms having a p-value below the critical value (0.0125) were considered significant. The non-significant terms below the double line did not discriminate the products.
Fig. 4 (top row) shows a heatmap visualisation of the univariate (Test 2) p-values in Table 2. Terms that significantly discriminate the products are shown in green, with deeper colours indicating a more significant test result. Terms above the BH critical value, which are non-significant across the products, are shown in light grey.
@@
Cluster analysis of the terms, as described in Section 2.5, organises the terms into a two-cluster solution, shown at the top of Fig. 4. Positively valenced terms (cluster 1) are contrasted against negatively valenced terms (cluster 2). Term cluster 1 consists of 15 positively valenced terms, 11 of which discriminated the products: Comforted, Guilty pleasure, Trust, Approval, Pleasant surprise, Interested, At ease, Pleased, Satisfaction, Happy and Good. The lexicon developers did not classify the term Guilty pleasure as either positively or negatively valenced (Ng et al., 2013a), but these term clusters suggest that consumers conceptualised Guilty pleasure as positively valenced. The remaining four terms in this cluster (Desire, Warm, Attentive, Reminiscence) did not discriminate the products. Term cluster 2 is comprised of 19 negatively valenced terms, of which 11 discriminated the products significantly (Disappointment, Displeasure, Annoyed, Shocked, Discontent, Uncomfortable, Disgust, Unhappy, Unpleasant surprise, Sickly, Regret) and eight did not (Bored, Curious, Angry, Worried, Cautious, Confused, Sceptical, Resentment).
Fig. 4 (lower part) also shows a $P\times T$ heatmap visualization of the elementwise (Test 3) p-values obtained by analysing, for each term, which of the $P$ products differs from the median value. We decided a priori to conduct all $P\times T=11\times 34=374$ elementwise tests (Section 2.3) with the FDR controlled at level 0.05 (Section 2.4). In 374 elementwise tests, 89 p-values were less than or equal to the BH critical value of $0.0118$.
@@
At least one product differed significantly from the median for all but four terms: Attentive, Worried, Cautious, and Confused. These four terms are among the 12 non-significant terms (coloured light grey in the top row of Fig. 4). Inspection of the columns for the other eight non-significant terms show at least one significant elementwise test. This result shows that individual products can differ significantly from the median of a term, even if the product citation percentages do not differ significantly overall on that term.
@@
Fig. 4 shows the dendrogram from the complete-linkage cluster analysis of products to the left of the elementwise heatmap. The rows (products) are sorted to accommodate this dendrogram which shows a crisp two-cluster solution. Product cluster 1 is composed of five standard products with added sugar (P5sw, P7sw, P11sw, P1sw, P9sw). Product cluster 2 is composed of six more varied products: four products without added sugar (P10eo, P8so, P2so and P6so) and two niche products with added sugar (P3nw and P4nw).
@@
The elementwise heatmap reveals approximate characterisations of products in the two product clusters. Members of product cluster 1 elicited significantly higher citation percentages than the median product in positively valenced terms in term cluster 1 and significantly lower citation percentages than the median product in negatively valenced terms in term cluster 2. By contrast, members of product cluster 2 elicited significantly higher citation percentages than the median product in negatively valenced terms in term cluster 2 and significantly lower citation percentages than the median product in positively valenced terms in term cluster 1.
@@
We draw attention to selected product results (rows) in the elementwise heatmap related to product cluster 2. Like other members of product cluster 2, the economical product P10eo had higher-than-median citation percentages for negatively valenced terms in term cluster 2 and lower-than-median citation percentages for positively valenced terms in term cluster 1. Additionally, P10eo elicited significantly higher-than-median citation percentages for Bored, Sceptical, and Resentment and a significantly lower-than-median citation percentage for Curious. Unlike the other 10 products, P10eo had relatively high intensities of catty, confectionery blackcurrant and artificial sweetness (Ng et al., 2012), which might explain why this product elicited these negatively valenced terms.
@@
Three standard products without added sugar (P2so, P6so and P8so) were also allocated to cluster 2. Besides of the general trend of product cluster 2, P6so and P8so elicited higher-than-median citation percentages for Sickly in term cluster 2. In a previous study, P6so and P8so were observed to have the highest intensities of artificial sweetness and natural processed blackcurrant (Ng et al., 2012), suggesting the relatively high Sickly citation rates were due to cloying oversweetness and processed flavour. These same sensory properties could be the reason for a lower-than-median citation percentage of Trust for P6so.